OpenStudy (anonymous):

Which of the following is equivalent to 3^5 = 243?

OpenStudy (anonymous):

log5 243 = 3

OpenStudy (anonymous):

I don't understand the question. Is something missing?

OpenStudy (anonymous):

no nothing is missing. its asking what log it is

OpenStudy (anonymous):

log3243 = 5 log5243 = 3 log35 = 243 log53 = 243

OpenStudy (anonymous):

There is a kind of rule / thing with the logarithm and exponential functions. It goes like this: log a x = y iff a^y = x. So, if you put the stuff above in, for your case ( 3^5 = 243 ) you'll see that a=3, y=5 and x = 243. Then you can use the inverse thing, which says log 3 243 = 5.

OpenStudy (anonymous):

it is multiple choice? which one is correct?

OpenStudy (anonymous):

i think its 5 234 = 5

OpenStudy (anonymous):

243*

OpenStudy (anonymous):

rewrite in equivalent exponential form and see which one is right \[\log_b(x)=y\iff b^y=x\]

OpenStudy (anonymous):

cus dont u put the base back where it should be..idk i remember something bout changin things around but the numbers dont change.

OpenStudy (anonymous):

\[log_3(243) = 5\iff 3^5=243\]

OpenStudy (anonymous):

ok so i was right

OpenStudy (anonymous):

yes that is true!

OpenStudy (anonymous):

log5 25 would be log 5 over log 25

OpenStudy (anonymous):

you should be able to switch back and forth easily between \[b^x=y\] and \[\log_b(y)=x\] it will make your life easier (at least in math class)

OpenStudy (anonymous):

\[\log_5(25)=y\iff 5^y=25\]

OpenStudy (anonymous):

therefore \[\log_5(25)=2\] because \[5^2=25\]

OpenStudy (anonymous):

i know but its asking what it would be for the change of base formula

OpenStudy (anonymous):

so like fraction form

OpenStudy (anonymous):

From wikipedia: \[ \log_b(x) = \frac{\log_k(x)}{\log_k(b)}\]

OpenStudy (anonymous):

log b(x) is the form you currently have it in, and if you want to calculate it in some base k, then you do that by the formula given by wikipedia, on the right, up there.